# Molecular Fingerprints Are Strong Models for Peptide Function Prediction

### v1

@@ Abstract
We study the effectiveness of molecular fingerprints for peptide property prediction and demonstrate that domain-specific feature extraction from molecular graphs can outperform complex and computationally expensive models such as GNNs, pretrained sequence-based transformers and multimodal ensembles, even without hyperparameter tuning. To this end, we perform a thorough evaluation on 126 datasets, achieving state-of-the-art results on LRGB and 5 other peptide function prediction benchmarks. We show that models based on count variants of ECFP, Topological Torsion, and RDKit molecular fingerprints and LightGBM as classification head are remarkably robust. The strong performance of molecular fingerprints, which are intrinsically very short-range feature encoders, challenges the presumed importance of long-range interactions in peptides. Our conclusion is that the use of molecular fingerprints for larger molecules, such as peptides, can be a computationally feasible, low-parameter, and versatile alternative to sophisticated deep learning models.
Machine Learning, ICML
@@ 1 Introduction
Peptides are small proteins, typically with amino acid chains of 3333-$5050$ residues and hundreds of atoms (Singh et al., 2015). They perform a variety of biological functions and have attracted increasing attention from researchers, mainly in the context of peptide-based drugs. Their useful therapeutic properties include antiviral, antitoxin, and combating cancer cells (Goles et al., 2024). They are also exceptionally effective in treating obesity and diabetes (Wang et al., 2022). In particular, antimicrobial peptides (AMPs) offer a potential solution to the escalating antimicrobial resistance crisis by offering an alternative to traditional antibiotics (Szymczak et al., 2023; Santos-Júnior et al., 2024).
Peptide function prediction problems are not only important in practice, but also manifest specific challenges in building machine learning pipelines, such as negative class sampling (Sidorczuk et al., 2022) or homology bias (Fernández-Díaz et al., 2024). This motivated the creation of multiple benchmark datasets that promote fair comparison with predefined splits, varied targets, and established evaluation protocols.
Peptides can be represented in many different modalities, with the selected representation determining the applicable machine learning models. When stable spatial folding structure is available, structural encodings can be used (Spänig et al., 2021), but this is relatively rare for peptides. In bioinformatics, typical approaches rely on complex feature engineering in amino acid sequences (Sidorczuk et al., 2022), with newer methods leveraging pretrained Protein Language Models (PLMs) (Elnaggar et al., 2022). Alternatively, peptides can be represented at the atom level by utilizing their topological molecular graphs (without modeling the spatial structure). This approach has been particularly popular in the machine learning literature due to the rapid development of Graph Neural Networks (GNNs) and Graph Transformers (GTs). Long Range Graph Benchmark (LRGB) (Dwivedi et al., 2022) in particular proposed the peptide function prediction as a graph-level, long-range interaction problem. It has been extensively used to benchmark the ability of novel GNN architectures to effectively capture long range dependencies between graph nodes.
In chemoinformatics, molecular fingerprints (Todeschini & Consonni, 2009) are one of the most popular models for feature extraction from small molecules, widely applied to similarity search, chemical space exploration, and molecular property prediction (Muegge & Mukherjee, 2016). They have also shown remarkable performance compared to GNNs (Jiang et al., 2021; Baptista et al., 2022; Xia et al., 2023), indicating the importance of domain-specific encoding algorithms. However, the adoption of this approach to predict peptide functions remains very limited, with just a few hybrid models that incorporate binary fingerprints as one component of their approach (Kang et al., 2024; Tan et al., 2024; Liang et al., 2024). Recent studies (Jiang et al., 2021; Tönshoff et al., 2024; Adamczyk & Czech, 2024) suggest that domain-specific features, paired with a robust nonlinear classifier, can surpass complex models in performance on both small and large molecules, achieving impressive results also on the LRGB benchmark.
In this study, we revisit the effectiveness of count-based variants of hashed molecular fingerprints for predicting peptide functions. These models, by design, are highly localized, counting occurrences of small, short-range subgraphs. When combined with a LightGBM classifier, fingerprint representations achieve state-of-the-art performance on the LRGB benchmark as well as five additional peptide function prediction benchmarks, spanning a total of 126 datasets.
The key contributions of this work are as follows. First, we show that count-based variants of hashed molecular fingerprints, coupled with the LightGBM classifier, obtain SOTA results in varied peptide function prediction tasks, with other advantages, such as simplicity and computational efficiency. We benchmark this method on the largest (to our knowledge) collection of 6 benchmarks and 126 datasets. Secondly, we show that these inherently very short-range models outperform long-range GNNs, challenging the presumed importance of long-range interactions in peptides and their viability for benchmarking long-range graph models.
@@ 2 Related works
The wide variety of prediction models for peptide functions arises from differences in input representations (e.g. sequence-based vs. graph-based, structural encodings, or feature extraction vs. learned representations) and reflects the influence of distinct research communities and the diversity of underlying datasets.
Considering granular, atom-level peptide representation in a form of molecular graph, Graph Neural Networks (GNNs) provide an automated method for building prediction models capable of incorporating not only structural information but also atom and bond attributes. The field of graph representation learning has advanced significantly since the introduction of foundational models such as Graph Convolutional Network (GCN) (Kipf & Welling, 2017), GraphSAGE (Hamilton et al., 2017) or Graph Isomorphism Network (GIN) (Xu et al., 2019), paving the way for more expressive models such as GINE (Hu et al., 2020) or GatedGCN (Dwivedi et al., 2023). However, these GNNs, limited by their reliance on iterative one-hop message passing, struggle to effectively capture information from distant nodes, which can be crucial for classification of larger molecules. In contrast, Graph Transformers (GT) overcome this limitation by allowing direct attention between all nodes, yet their lack of inherent graph inductive biases forces them to depend on arbitrary positional encodings, undermining their ability to fully exploit graph structure. The tradeoff between those two factors led to the development of modern hybrid GNN architectures, which display much better accuracy on molecular benchmarks. GraphGPS (Rampášek et al., 2022) combines local message passing with global attention and positional encodings, helping the attention mechanism understand the relative positions of the nodes. Enhanced using the Hierarchical Distance Structural Encoding (HDSE) mechanism (Luo et al., 2024), GraphGPS achieves one of the best results in the Peptide-func dataset from LRGB. In (Gutteridge et al., 2023), the Dynamically Rewired Message Passing (DRew) model was introduced. It is an adaptable framework, for any message-passing GNN, incorporating layer-dependent rewiring, controlled message passing delays, and skipping connections. On the other hand, the graph inductive bias transformer (GRIT) (Ma et al., 2023) does not require message passing, but still captures local and global graph properties using degree information injection and learned relative positional encodings. Subsequently, other approaches have been proposed to capture long-range dependencies. Graph ViT (He et al., 2023) is a unique model for graph representation learning that integrates the strengths of GNN-based encoders for local feature extraction with Vision Transformers to capture global context effectively. Finally, S2GNNs (Spatio-Spectral Graph Neural Networks) (Geisler et al., 2024) are a novel type of graph neural network that combines spatial message passing with spectral graph filters and have recently been shown to outperform state-of-the-art graph transformers in peptide function prediction tasks.
Due to the limited data availability, many peptide prediction tasks involve only around a thousand samples or fewer. Furthermore, simple features like amino acid compositions are known to be strong predictors of many properties. As a result, the use of deep learning remains limited in bioinformatics tasks on peptides. Instead, numerous feature extraction methods, often quite sophisticated, have been developed. They often utilize multiple modalities, extracting physicochemical properties of peptide molecules (e.g. amphiphilicity, hydrophobicity), as well as compositional features from amino acid sequences. Such methods include MLAMP (Lin & Xu, 2016), ampir (Fingerhut et al., 2020), and MACREL (Santos-Júnior et al., 2020) or amPEPpy (Lawrence et al., 2021). Combined with traditional classifiers such as SVM and Random Forest, they have shown good performance, even compared to deep learning methods like AMPScannerV2 (Veltri et al., 2018).
Novel deep learning models on amino acid sequences are based on pretrained transformers. Those Protein Language Models (PLMs) are trained on large-scale datasets with proteins typically much larger than peptides. They are heavily inspired by natural language processing (NLP) and include, e.g. ProtBERT and Prot-T5-XL (Elnaggar et al., 2022). The authors of the ESM architecture (Lin et al., 2023) showed that they exhibit a scaling behavior similar to large language models (LLMs) in NLP. ESM models have also shown particularly strong performance in independent benchmarks compared to other PLMs (Fernández-Díaz et al., 2024).
Since antimicrobial peptides (AMPs) are particularly important in drug design, there are many specialized PLMs for predicting this property, including BERT-Protein (Zhang et al., 2021), LMPred (Dee, 2022), cAMPs_pred (Ma et al., 2022), and AMP-BERT (Lee et al., 2023). They can be used out-of-the-box with pretrained AMP classification heads.
@@ 3.1 Molecular fingerprints
Molecular fingerprints are feature extraction algorithms for molecules. The vast majority of them are based on extracting small subgraphs and detecting their presence or counting occurrences (binary and count variants, respectively). This turns the problem of molecular graph classification into tabular classification, typically combined with tree-based ensemble classifiers such as Random Forest (Breiman, 2001) or LightGBM (Ke et al., 2017).
@@
We focus on hashed fingerprints (Todeschini & Consonni, 2009), exemplified by ECFP (Rogers & Hahn, 2010) or Topological Torsion (Nilakantan et al., 1987). They are more flexible than structural fingerprints like MACCS or PubChem, which use predefined subgraphs like functional groups or rings of given size, and are not designed for peptides. Hashed fingerprints use a more flexible approach, defining a general “shape” of extracted subgraphs, for example, circular atom neighborhoods in ECFP (Rogers & Hahn, 2010), short paths in Topological Torsion (TT) (Nilakantan et al., 1987), or all small subgraphs in RDKit fingerprint (rdk, ). Extracting subgraph features in this manner offers greater flexibility, as it does not, for instance, rely on specific atom types. Subgraphs are defined by their structure, including the specific atoms and bonds they contain (e.g. characterized by atomic numbers and bond types), which are then converted into a unique integer identifier. To obtain a constant-length feature vector, they are hashed into the output vector, using the modulo function in the simplest case. The binary variant simply indicates whether a given substructure appears in a molecule at all, while the count-based variant tallies the occurrences, also accounting for potential hashing collisions.
@@
ECFP typically uses circular subgraphs of radius 2 (see Figure 1). It starts with the atom itself, creating a numerical identifier based on atomic number, number of heavy neighbors, valence, and a few other simple features (see (Rogers & Hahn, 2010) for a complete list). Then, it iteratively increases the radius, incorporating neighbors by combining their identifiers, up to a given radius. This results in the final subgraph identifier for each atom. Such a procedure, known as the Morgan algorithm, closely resembles the operation of message passing in GNNs, as both are rooted in the Weisfeiler-Lehman isomorphism test. As such, ECFP with radius 2 is roughly equivalent to a shallow 2-layer GNN. Topological Torsion (Nilakantan et al., 1987), designed to model short-range molecular interactions, encodes linear paths of length 4, combining the identifiers of these consecutive atoms. RDKit fingerprint (rdk, ) uses all subgraphs of size up to 7 bonds, which can be nonlinear and include cycles (e.g. rings). The major difference between TT and RDKit is that the former uses only linear paths, while the RDKit fingerprint also encodes small cyclic structures (e.g. benzene rings).
@@
We focus on those three fingerprints, as they are strictly local, very short-range descriptors. Therefore, a good performance of models based on their features can be attributed only to the importance of the short-range relations, rather than to long-range dependencies, since they are inherently incapable of capturing them. This locality makes them notably fast to compute (Adamczyk & Ludynia, 2024) and scalable to larger molecules like peptides, e.g. ECFP scales linearly with the number of atoms. Because peptides are relatively linear structures, path-based TT and RDKit fingerprints also scale well. Those three fingerprints have also shown remarkable performance in ligand-based virtual screening (Riniker & Landrum, 2013).
@@
Figure 1: Example subgraph extraction in consecutive iterations of ECFP algorithm. Reprinted with permission from (Rogers & Hahn, 2010). Copyright 2010 American Chemical Society.
@@ 3.2 Long-range interactions (LRI)
Long Range Graph Benchmark (LRGB) (Dwivedi et al., 2022) proposed 3 conditions for a dataset to potentially require long-range interactions (LRI). Firstly, a necessary condition is that the graphs have to be sufficiently large, in particular larger than typical GNNs receptive field. Secondly, the nature of the task itself has to indicate that long-range dependencies between nodes are crucial to properly predict properties, with local interactions not being enough. Lastly, the experimental results should support this, i.e. models that incorporate global knowledge, like GNNs with topological encodings or graph transformers, have to outperform local message-passing ones. For graph classification and regression, peptide property prediction was proposed as a group of tasks that meet these conditions.
We show that this is not necessarily the case, particularly in relation to the third assumption. Peptides at the atomic level are certainly large and “long” (large diameter, low atom degrees) enough graphs to suggest a possible LRI. However, the biological motivation for LRI in peptides is somewhat arguable. Proteins have long-range dependencies because of their large structures, which enable complex folding through distant amino acid interactions. In contrast, the smaller size of the peptides naturally limits such possibilities.
The third point, i.e. the superior performance of models with global knowledge, is the easiest to validate quantitatively. Since the argument is based on experimental results, it suffices to show that inherently local short-range models outperform long-range ones. Molecular fingerprints exactly fulfill these requirements. By obtaining SOTA results, we show that they provide strong evidence against the third claim.
@@ 4 Experiments and results
We train proposed fingerprint-based models on a variety of benchmarks that cover a wide range of peptide properties, data sampling methods, quality metrics, and train-test splits. We adhere closely to the procedure outlined for each benchmark in its original publication and report the corresponding quality metrics. For benchmarks that provide multiple metrics, we present AUROC and Matthews Correlation Coefficient (MCC) here due to space constraints, with the remaining metrics detailed in Appendix A. The best result for each table is highlighted in bold.
Our proposed fingerprint-based models always use their count variants and are implemented with the scikit-fingerprints (Adamczyk & Ludynia, 2024) library. We use default values for all hyperparameters: ECFP radius 2 (diameter 4, also known as ECFP4), Topological Torsion path length 4, and RDKit fingerprint path length 7. They use small subgraphs as features, capturing only very short-range dependencies. We did not see any significant improvement when tuning those hyperparameters (except for PeptideReactor, see Section 4.3).
We use LightGBM (Ke et al., 2017) with 500 trees as the classification model, due to its robust performance. Since most datasets are heavily imbalanced, we use class weight balancing inversely proportional to the frequency of the positive class. We keep default values for all other hyperparameters because we noticed negligible improvements when tuning them. Only setting the number of trees higher than the default 100 was relevant for performance, which is consistent with the literature (Probst et al., 2019).

### v2

@@ Abstract
Motivation: Understanding peptide properties is often assumed to require modeling long-range molecular interactions, motivating the use of complex graph neural networks and pretrained transformers. Yet, whether such long-range dependencies are essential remains unclear. We investigate if simple, domain-specific molecular fingerprints can capture peptide function without these assumptions. Atomic-level representation aims to provide richer information than purely sequence-based models and better efficiency than structural ones.
Results: Across 132 datasets, including LRGB and five other peptide benchmarks, models using count-based ECFP, Topological Torsion, and RDKit fingerprints with LightGBM achieve state-of-the-art accuracy. Despite encoding only short-range molecular features, these models outperform GNNs and transformer-based approaches. Control experiments with sequence shuffling and amino acid counts confirm that fingerprints, though inherently local, suffice for robust peptide property prediction. Our results challenge the presumed necessity of long-range interaction modeling and highlight molecular fingerprints as efficient, interpretable, and computationally lightweight alternatives for peptide prediction.
Contact: jadamczy@agh.edu.pl
Supplementary information: All code and data are available at:
https://github.com/scikit-fingerprints/peptides_molecular_fingerprints_classification
keywords:
peptide function prediction, machine learning, chemoinformatics, molecular fingerprints
@@ 1 Introduction
Peptides are small proteins, typically composed of 3–50 amino acids, that perform diverse biological functions and are increasingly being studied as potential therapeutics. They can bind to proteins, regulate their activity and show antiviral, antitoxin, anticancer, and antidiabetic properties peptides_drug_discovery ; peptides_applications_diabetes . In particular, antimicrobial peptides (AMPs) are promising candidates against the growing antimicrobial resistance crisis peptides_applications_amp . Therefore, accurate prediction of peptide properties is essential for drug discovery.
Peptide function prediction poses unique challenges for machine learning (ML), including homology bias and imbalanced datasets AutoPeptideML ; AMPBenchmark . Consequently, numerous benchmark datasets with standardized evaluation procedures have been created. Peptides can be represented through various modalities, influencing the choice of ML models. When 3D structures are available, structural encodings can be used PeptideReactor , but these are rare and computationally expensive to obtain. Thus, sequence- and graph-based representations that avoid folding simulations have gained more attention.
Recent years have seen a shift from handcrafted sequence features to pretrained protein language models (PLMs) such as ProtBERT and ESM ProtTrans ; ESM2 . Alternatively, peptides can be modeled as molecular graphs, enabling the use of Graph Neural Networks (GNNs) and Graph Transformers (GTs). The Long-Range Graph Benchmark (LRGB) LRGB formulated peptide prediction as a long-range dependency problem. However, the practical importance of long-range interactions in short and flexible peptides remains unclear.
In chemoinformatics, molecular fingerprints - compact encodings derived from small molecular subgraphs molecular_descriptors - are widely used for prediction of small-molecule properties. They offer domain-specific, efficient representations that often rival deep learning methods fingerprints_vs_gnns ; molecular_property_predictions_gnn_limitations . Despite their success, their use for peptide prediction has been limited, with only a few hybrid approaches that incorporate binary fingerprints MMDB ; peptide_permeability_prediction .
In this work, we revisit count-based molecular fingerprints as features for peptide function prediction. When combined with a LightGBM classifier, these representations achieve state-of-the-art results across six benchmarks, including LRGB, without requiring hyperparameter tuning or structural information. This suggests that short-range subgraph statistics may be sufficient to capture key biochemical features, challenging the assumption that long-range dependencies are essential.
Our contributions are threefold: (1) we demonstrate that count-based molecular fingerprints combined with LightGBM outperform complex deep learning models on 132 datasets, (2) we show that these localized descriptors can match or surpass long-range GNNs, and (3) we provide the most comprehensive benchmark of fingerprint-based peptide models to date.
@@ 2 Related works
Prediction approaches differ mainly by representation. For atom‑level peptide graphs, message‑passing GNNs such as GCN, GraphSAGE and GIN learn local structure, but have limited receptive fields GCN ; GraphSAGE ; GIN . Graph Transformers use global attention, yet require positional encodings and have weak graph inductive bias. Hybrid or alternative designs, such as GraphGPS (with HDSE), GRIT, GraphViT, and S2GCN, seek to capture both local and long‑range structure and report strong results on molecular benchmarks including LRGB peptides dataset rampavsek2022recipe ; HDSE ; GRIT ; GraphViT ; S2GCN .
Sequence-based bioinformatics pipelines remain competitive, especially when data is scarce. They derive physicochemical and compositional descriptors (e.g. amphiphilicity, hydrophobicity) and use classical classifiers like RandomForest or SVM. Notable examples include Ampir, MACREL and amPEPpy ampir ; MACREL ; lawrence2021ampeppy , which can rival early deep models such as AMPScannerV2 veltri2018deep . More recently, pretrained transformers (PLMs) such as ProtBERT and ProtT5 gained popularity. Those models are pretrained on large protein corpora, including peptide sequences ProtTrans . The ESM family of models shows a scaling behavior similar to LLMs and give strong results on independent peptide benchmarks ESM2 ; AutoPeptideML . Focusing on AMPs, specialized PLMs, e.g., BERT‑Protein, LMPred, cAMPs‑pred, and AMP‑BERT, are available with pretrained heads for direct use BERT_Protein ; dee2022lmpred ; ma2022identification ; lee2023amp .
Our work differs in adopting domain‑specific, count-based molecular fingerprints as peptide structure encoders. Compared with deep models, fingerprints provide robust, low‑parameter features and, as we show, compete or excel in peptide tasks without requiring long‑range mechanisms.
@@ 3.1 Molecular fingerprints
Molecular fingerprints are feature extraction algorithms for molecules. The vast majority of them are based on extracting small subgraphs and detecting their presence or counting occurrences (binary and count variants, respectively). This turns the problem of molecular graph classification into tabular classification, typically combined with tree-based ensemble classifiers such as Random Forest or LightGBM LightGBM . See Figure 1 for visualization of the proposed pipeline.
Figure 1: The proposed pipeline for peptide function prediction.
@@
We focus on hashed fingerprints molecular_descriptors , exemplified by ECFP ECFP or Topological Torsion topological_torsion . They are more flexible than structural fingerprints like MACCS or PubChem, which use predefined subgraphs like functional groups or rings of given size, and are not designed for peptides. Hashed fingerprints use a more flexible approach, defining a general “shape” of extracted subgraphs, for example, circular atom neighborhoods in ECFP ECFP , paths of given length in Topological Torsion (TT) topological_torsion , or all small subgraphs in RDKit fingerprint rdkit_fp . Extracting subgraph features in this manner offers greater flexibility, as it does not, for instance, rely on specific atom types. Subgraphs are defined by their structure, including the specific atoms and bonds they contain (e.g. characterized by atomic numbers and bond types), which are then converted into a unique integer identifier. To obtain a constant-length representation, they are hashed into the output vector, using the modulo function in the simplest case. The binary variant only indicates whether a given substructure appears in a molecule at all, while the count-based variant tallies the occurrences, incorporating more information about the compound structure and size.
@@
ECFP typically uses circular subgraphs of radius 2 (see Figure 2). It starts with the atom itself, creating a numerical identifier based on atomic number, number of heavy neighbors, valence, and a few other simple features (see ECFP for a complete list). Then, it iteratively increases the radius, incorporating neighbors by combining their identifiers, up to a given maximal radius. This results in the final subgraph identifier for each atom. Such a procedure, known as the Morgan algorithm, closely resembles the operation of message passing in GNNs, as both are rooted in the Weisfeiler-Lehman isomorphism test. Topological Torsion topological_torsion , designed to model short-range molecular interactions, encodes linear paths of length 4, combining the identifiers of these consecutive atoms. RDKit fingerprint rdkit_fp uses all subgraphs of size up to 7 bonds, which can be nonlinear and include cycles (e.g. rings). The major difference between TT and RDKit is that the former uses only linear paths, while the RDKit fingerprint also encodes small cyclic structures (e.g. benzene rings).
The shapes of subgraphs extracted by ECFP with radius 2 are roughly equivalent to those of a shallow 2-layer GNN GNNs_as_continous_fingerprints . However, fingerprints do not learn task-specific weights like GNNs, instead performing domain-specific, deterministic feature extraction. This naturally reduces their risk of overfitting and increases the robustness of the learned feature space fingerprints_vs_gnns ; molecular_property_predictions_gnn_limitations . It also results in lower capacity of fingerprint-based models, which is beneficial in small, noisy datasets typical in chemo- and bioinformatics.
@@
We focus on the three aforementioned fingerprints, as they are strictly local, very short-range descriptors. Therefore, a good performance of models based on their features can be attributed only to the importance of the short-range relations, rather than to long-range dependencies, since they are inherently incapable of capturing them. This locality makes them notably fast to compute scikit_fingerprints and scalable to larger molecules like peptides, e.g. ECFP scales linearly with the number of atoms. Because peptides are relatively linear structures, path-based TT and RDKit fingerprints also scale well. Those three fingerprints have also shown remarkable performance in ligand-based virtual screening virtual_screening .
Lastly, we note that these molecular fingerprints do not require any information about molecular conformation, peptide folding, or even whether a stable structure exists at all. Atom-level graphs can be quickly and deterministically constructed from an amino acid sequence. See Figure 3 for an example, which can be directly vectorized by the molecular fingerprint.
@@
Figure 2: Example subgraph extraction in consecutive iterations of ECFP algorithm. Reprinted with permission from ECFP . Copyright 2010 American Chemical Society.
Figure 3: Atom-level molecular graph for Liraglutide peptide.
@@ 3.2 Long-range interactions (LRI)
Long Range Graph Benchmark (LRGB) LRGB proposed 3 conditions for a dataset to potentially require long-range interactions (LRI). Firstly, a necessary condition is that the graphs have to be sufficiently large, in particular larger than typical GNNs receptive field. Secondly, the nature of the task itself has to indicate that long-range dependencies between nodes are crucial to properly predict properties, with local interactions not being enough. Lastly, the experimental results should support this, i.e., models that incorporate global knowledge, like GNNs with topological encodings or graph transformers, have to outperform local message-passing ones. For graph classification and regression, peptide property prediction was proposed as a group of tasks that meet these conditions.
We show that this is not necessarily the case, particularly in relation to the third condition. Peptides at the atomic level are certainly large and “long” (large diameter, low atom degrees) enough graphs to suggest a possible LRI. However, the biological motivation for LRI in peptide function prediction is weaker compared to larger proteins. Long-range interactions are relevant in proteins because of their large structures, which enable complex folding through distant amino acid interactions, and thus impact their functional properties. In contrast, the smaller size of the peptides naturally limits such possibilities, in many cases leading to a lack of well-defined structure peptides_structure .
The third point, i.e., the superior performance of models with global knowledge, is the easiest to validate quantitatively. Since the argument is based on experimental results, it suffices to show that inherently local short-range models outperform long-range ones. Molecular fingerprints exactly fulfill these requirements. By obtaining SOTA results, we show that they provide strong evidence against the third claim, i.e., that peptide function prediction relies primarily on short-range interactions.
@@ 4 Experiments and results
We train the proposed fingerprint-based models on a variety of benchmarks that cover a wide range of peptide properties, data sampling methods, quality metrics, and train-test splits. We adhere closely to the procedure outlined for each benchmark in its original publication and report the corresponding quality metrics. For benchmarks that provide multiple metrics, we present AUROC and Matthews Correlation Coefficient (MCC) here due to space constraints; with the remaining metrics detailed in the Supplementary Material. The best result for each table is highlighted in bold.
Our proposed fingerprint-based models always use their count variants and are implemented with the scikit-fingerprints scikit_fingerprints library. We use default values for all hyperparameters: ECFP radius 2 (diameter 4, also known as ECFP4), Topological Torsion path length 4, and RDKit fingerprint path length 7. They use small subgraphs as features, capturing only very short-range dependencies. We did not see any significant improvement when tuning those hyperparameters (except for PeptideReactor, see Section 4.3).
We use LightGBM LightGBM with 500 trees as the classification model, due to its robust performance. Since most datasets are heavily imbalanced, we use class weight balancing inversely proportional to the frequency of the positive class. We keep default values for all other hyperparameters because we noticed negligible improvements when tuning them. Only setting the number of trees higher than the default 100 was relevant for performance, which is consistent with the literature tunability_probst .
The low sensitivity of the proposed method to hyperparameter values is an advantage as it reduces the computational cost for new tasks, particularly for large-scale virtual screening of potential peptide therapeutics. It also makes it a fast and easy-to-use baseline for other work in this area.